hello- which set of points are on the line y=2x-1?
Answer by rothauserc(4718)

we note that y = 2x -1 is already in the general slope / y intercept form y = mx +b where m is the slope and b is the y intercept
so, m = 2 and the point (0, -1) is the y intercept
m = 2 means that to find a point on the line we just move 2 grid points up (y value) and 1 grid point over (x value) from a point on the line, i.e. the y intercept, so here we go
set of points on the line y=2x-1 = (0, -1), (1, 1), (2, 3), (3, 5), (4, 7) and so forth......
you can also graph the line by using the points